OpenStudy (anonymous):

Given two spheres, if the volume of the first sphere is 4/3*3.14 and the volume of the second sphere is 36*3.14 , which is the relation of the second radius to the first? squared tripled doubled multiplied by a factor of 8/3

OpenStudy (anonymous):

the volume of the sphere v=(4/3)pi*r^3 v1=(4/3)pi, here r1=1 v2=36pi v2=36pi=(4/3)pi*r^3 36=(4/3)*r^3 36/(4/3)=r^3 27=r^3 cube root of(27)=r 3sqrt3=r=r2 r2 is about three times the r1 ..tripled ans